Please help me solve this solution, QUESTION: In a certain school it was found that 28% of the children came to school by car. The number of children coming by car was 231. There were 517 children walking to school. The other children came by PMV. What percentage of the children came by PMV?
Answer by thesvw(77)

Lets find the total number of students. First understand what 28% really say. It says if theres hundred students in school, 28 come by car.
Total number of students = 231x 100/28 = 825
517 walks. So the remaining goes by PMV, number of students go by PMV = 308
So the percentage = 517/825 x 100 = 62.67%
Dont forget to thank.